11123:

At a conference for 500 people, 10% of the participants are French, 35% are Americans, 20% are Germans, and are of other nationalities. How many French participants are at the conference? 10 20 40 50 Question 10(Multiple Choice Worth 5 points) (07.01 MC) Peter is giving marbles to some children at a carnival. He has 5 red marbles, 4 blue marbles, and 3 yellow marbles. If Peter selects a marble randomly without looking, what is the probability that he will give a blue marble to the first child and then a yellow marble to the second child? 4 over 12 plus 3 over 12 is equal to 7 over 12 4 over 12 plus 3 over 11 is equal to 80 over 132 4 over 12 multiplied by 3 over 12 is equal to 12 over 144 4 over 12 multiplied by 3 over 11 is equal to 12 over 132 Question 11(Multiple Choice Worth 5 points) (07.03 MC) Kane used a probability simulator to roll a 6-sided number cube and flip a coin 100 times. The results are shown in the tables below: Number on the Cube Number of Times Rolled 1 18 2 32 3 12 4 25 5 3 6 10 Heads Tails 55 45 Using Kane's simulation, what is the probability of rolling a 2 on the number cube and the coin landing heads up? 1760 over 10000 1620 over 10000 87 over 100 67 over 100 Question 12(Multiple Choice Worth 5 points) (07.01 LC) Abel makes the following statement: "I have a 30% chance of being selected as the captain of the baseball team and a 20% chance of being elected student body president." What is the probability that Abel will be selected as the captain of the baseball team and will be elected student body president? 6% 10% 20% 50%

moreso:

500 people, 10% are french multiply 10/100 * 500

11123:

so 50 ?

moreso:

yes

11123:

i mainly need help with 11

moreso:

Let me organize the data first Is this correct? Number on the Cube Number of Times Rolled 1 18 2 32 3 12 4 25 5 3 6 10 Heads 55 Tails 45

moreso:

probabilty ( roll 2 & heads) = 32/100 * 55/100 = 1760 / 10000

11123:

thx